\begin{thebibliography}{10}

\bibitem{ACKERMANN1973}
F.~Ackermann, I.~Platz, and G.~zu~Putlitz.
\newblock Nuclear quadrupole moments of radioactive83rb,84rb and86rb.
\newblock {\em Zeitschrift für Physik A Hadrons and nuclei 1973 260:2},
  260:87--110, 4 1973.

\bibitem{Ahmad1983}
S.~A. Ahmad, W.~Klempt, R.~Neugart, E.~W. Otten, K.~Wendt, and C.~Ekström.
\newblock Determination of nuclear spins and moments in a series of radium
  isotopes.
\newblock {\em Physics Letters B}, 133:47--52, 12 1983.

\bibitem{Allegrini22}
M.~Allegrini, E.~Arimondo, and Luis~A. Orozco.
\newblock Survey of hyperfine structure measurements in alkali atoms.
\newblock {\em Journal of Physical and Chemical Reference Data}, xx(x):xxxx,
  2022.

\bibitem{Armstrong1965}
R.~Marrus~L. Armstrong.
\newblock Magnetic moments of rhenium-186 and rhenium-188 and analysis of the
  rhenium hyperfine structure.
\newblock {\em Phys. Rev.}, 138:B310, 1965.

\bibitem{Arnold1987}
E.~Arnold, W.~Borchers, M.~Carre, H.~T. Duong, P.~Juncar, J.~Lerme,
  S.~Liberman, W.~Neu, R.~Neugart, E.~W. Otten, M.~Pellarin, J.~Pinard, G.~Ulm,
  J.~L. Vialle, and K.~Wendt.
\newblock Direct measurement of nuclear magnetic moments of radium isotopes.
\newblock {\em Physical Review Letters}, 59:771, 8 1987.

\bibitem{Atkinson2019}
Peggy~E. Atkinson, Jesse~S. Schelfhout, and John~J. McFerran.
\newblock Hyperfine constants and line separations for the <math xmlns.
\newblock {\em Physical Review A}, 100:042505, 10 2019.

\bibitem{Barzakh2020}
A.~E. Barzakh, D.~Atanasov, A.~N. Andreyev, M.~Al Monthery, N.~A. Althubiti,
  B.~Andel, S.~Antalic, K.~Blaum, T.~E. Cocolios, J.~G. Cubiss, P.~Van Duppen,
  T.~Day Goodacre, A.~De Roubin, Yu~A. Demidov, G.~J. Farooq-Smith, D.~V.
  Fedorov, V.~N. Fedosseev, D.~A. Fink, L.~P. Gaffney, L.~Ghys, R.~D. Harding,
  D.~T. Joss, F.~Herfurth, M.~Huyse, N.~Imai, M.~G. Kozlov, S.~Kreim,
  D.~Lunney, K.~M. Lynch, V.~Manea, B.~A. Marsh, Y.~Martinez Palenzuela, P.~L.
  Molkanov, D.~Neidherr, R.~D. Page, M.~Rosenbusch, R.~E. Rossel, S.~Rothe,
  L.~Schweikhard, M.~D. Seliverstov, S.~Sels, C.~Van Beveren, E.~Verstraelen,
  A.~Welker, F.~Wienholtz, R.~N. Wolf, and K.~Zuber.
\newblock Hyperfine anomaly in gold and magnetic moments of <math xmlns.
\newblock {\em Physical Review C}, 101:034308, 3 2020.

\bibitem{Barzakh2012}
A.~E. Barzakh, L.~Kh. Batist, D.~V. Fedorov, V.~S. Ivanov, K.~A. Mezilev, P.~L.
  Molkanov, F.~V. Moroz, S.~Yu. Orlov, V.~N. Panteleev, and Yu.~M. Volkov.
\newblock Hyperfine structure anomaly and magnetic moments of neutron deficient
  tl isomers with <span class.
\newblock {\em Physical Review C}, 86:014311, 7 2012.

\bibitem{Beckmann1974}
A.~Beckmann, K.~D. Böklen, and D.~Elke.
\newblock Precision measurements of the nuclear magnetic dipole moments
  of6li,7li,23na,39k and41k.
\newblock {\em Zeitschrift für Physik 1974 270:3}, 270:173--186, 9 1974.

\bibitem{BLACHMAN1969}
Arthur~G. Blachman, Donald~A. Landman, and Allen Lurio.
\newblock Hyperfine structure and <span class.
\newblock {\em Physical Review}, 181:70, 5 1969.

\bibitem{Bout1967}
Paul A.~Vanden Bout, Vernon~J. Ehlers, William~A. Nierenberg, and Howard~A.
  Shugart.
\newblock Hyperfine-structure separations, nuclear magnetic moments, and
  hyperfine-structure anomalies of gold-198 and gold-199.
\newblock {\em Physical Review}, 158:1078, 6 1967.

\bibitem{Brand1981}
H.~Brand, V.~Pfeufer, and A.~Steudel.
\newblock Laser-atomic-beam spectroscopy of 4f 75d6s - 4f 75d6p transitions in
  eu i.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1981 302:4},
  302:291--298, 12 1981.

\bibitem{BRASLAU1961}
Norman Braslau, Gilbert~O. Brink, and Jhan~M. Khan.
\newblock <span class.
\newblock {\em Physical Review}, 123:1801, 9 1961.

\bibitem{Brenner1985}
T.~Brenner, S.~Büttgenbach, W.~Rupprecht, and F.~Träber.
\newblock Nuclear moments of the low abundant natural isotope 176lu and
  hyperfine anomalies in the lutetium isotopes.
\newblock {\em Nuclear Physics A}, 440:407--423, 7 1985.

\bibitem{BROWN1966}
Howard~Howland Brown and John~G. King.
\newblock Hyperfine structure and octopole interaction in stable bromine
  isotopes.
\newblock {\em Physical Review}, 142:53, 2 1966.

\bibitem{BUDICK1970}
B.~Budick and J.~Snir.
\newblock Hyperfine-structure anomalies of stable ytterbium isotopes.
\newblock {\em Physical Review A}, 1:545, 3 1970.

\bibitem{Burt2009}
E.~A. Burt, S.~Taghavi-Larigani, and R.~L. Tjoelker.
\newblock High-resolution spectroscopy of <span class.
\newblock {\em Physical Review A}, 79:062506, 6 2009.

\bibitem{Burger1982}
K.~H. Bürger, B.~Burghardt, S.~Büttgenbach, R.~Harzer, H.~Hoeffgen,
  G.~Meisel, and F.~Träber.
\newblock Hyperfine structure and isotope shift of high-lying metastable states
  of rhenium.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei}, 307:201--209, 9
  1982.

\bibitem{Buttgenbach1977}
S.~Büttgenbach, R.~Dicke, H.~Gebauer, and M.~Herschel.
\newblock Hyperfine structure and nuclear moments of99ru and101ru.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1977 280:3},
  280:217--226, 9 1977.

\bibitem{Buttgenbach1978}
S.~Büttgenbach, R.~Dicke, H.~Gebauer, R.~Kühnen, and F.~Träber.
\newblock Hyperfine structure of six low-lying fine structure levels of 191ir
  and193ir and the 191 $\delta $ s 193 hyperfine anomaly.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1978 286:4},
  286:333--340, 12 1978.

\bibitem{Buttgenbach1981}
S.~Büttgenbach, R.~Dicke, G.~Gölz, and F.~Träber.
\newblock High precision hyperfine structure measurements in185re and187re.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1981 302:4},
  302:281--290, 12 1981.

\bibitem{Buttgenbach1974}
S.~Büttgenbach, M.~Herschel, G.~Meisel, E.~Schrödl, W.~Witte, and W.~J.
  Childs.
\newblock Ground state hyperfine structure of95mo and97mo.
\newblock {\em Zeitschrift für Physik 1974 266:4}, 266:271--274, 8 1974.

\bibitem{Chan1966}
Yau~W. Chan, Victor~W. Cohen, Max Lipsicas, and H.~B. Silsbee.
\newblock Nuclear magnetic moment and hyperfine-structure anomaly of <span
  class.
\newblock {\em Physical Review}, 150:933, 10 1966.

\bibitem{Chan1969}
Yau~W. Chan, Victor~W. Cohen, and H.~B. Silsbee.
\newblock Hyperfine structure and nuclear magnetic moment of <span class.
\newblock {\em Physical Review}, 184:1102, 8 1969.

\bibitem{CHANEY1969}
R.~L. Chaney and M.~N. McDermott.
\newblock Nuclear orientation of 113cdm, 115cd and 115cdm.
\newblock {\em Physics Letters A}, 29:103--104, 4 1969.

\bibitem{Chen2012}
Tzu-Ling Chen, Isaac Fan, Hsuan-Chen Chen, Chang-Yi Lin, Shih-En Chen,
  Jow-Tsong Shy, and Yi-Wei Liu.
\newblock Absolute frequency measurement of the <span class.
\newblock {\em Physical Review A}, 86:052524, 11 2012.

\bibitem{CHILDS1971}
W.~J. Childs.
\newblock Relativistic effects in the hyperfine structure of <span class.
\newblock {\em Physical Review A}, 4:439, 8 1971.

\bibitem{CHILDS1965}
W.~J. Childs and L.~S. Goodman.
\newblock Magnetic hyperfine structure of the <span class.
\newblock {\em Physical Review}, 137:A35, 1 1965.

\bibitem{CHILDS1979}
W.~J. Childs and L.~S. Goodman.
\newblock <span class.
\newblock {\em Physical Review A}, 20:1922, 11 1979.

\bibitem{CLARK1982}
D.~L. Clark and G.~W. Greenlees.
\newblock Changes of charge radii and hyperfine interactions of the dy
  isotopes.
\newblock {\em Physical Review C}, 26:1636, 10 1982.

\bibitem{Cochrane1998}
E~C~A Cochrane, D~M Benton, D~H Forest, and J~A~R Griffith.
\newblock Hyperfine structure and isotope shifts in natural vanadium.
\newblock {\em Journal of Physics B: Atomic, Molecular and Optical Physics},
  31:2203, 5 1998.

\bibitem{COHEN1962}
V.~W. Cohen, T.~Moran, and S.~Penselin.
\newblock Hyperfine structure of <span class.
\newblock {\em Physical Review}, 127:517, 7 1962.

\bibitem{CUSSENS1969}
C~J Cussens, G~K Rochester, and K~F Smith.
\newblock Measurements of the nuclear spin of 206tl, the hyperfine structure
  splitting of 66cu, and the nuclear magnetic moments of 108ag(2.3 min) and
  110ag(24s).
\newblock {\em Journal of Physics A: General Physics}, 2:658, 11 1969.

\bibitem{DAHMEN1967}
H.~Dahmen and S.~Penselin.
\newblock Measurement of the nuclear magnetic dipole moment of au197 and
  hyperfine structure measurements in the ground states of au197, ag107, ag109
  and k39.
\newblock {\em Zeitschrift für Physik 1967 200:4}, 200:456--466, 8 1967.

\bibitem{ECK1957}
T.~G. Eck, A.~Lurio, and P.~Kusch.
\newblock Hfs of the <span class.
\newblock {\em Physical Review}, 106:954, 6 1957.

\bibitem{EDER1985}
R.~Eder, E.~Hagn, and E.~Zech.
\newblock Nuclear magnetic resonance on oriented 4.7 h <span class.
\newblock {\em Physical Review C}, 32:1707, 11 1985.

\bibitem{EKSTROM1980}
C.~Ekström, L.Robertsson, S.~Ingelman, G.~Wannberg, and I.~Ragnarsson.
\newblock Nuclear ground-state spin of185au and magnetic moments of 187, 188au:
  Further evidence for coexisting nuclear shapes in this mass region.
\newblock {\em Nucl. Phys. A}, 348:25, 1980.

\bibitem{FAUST1960}
W.~Faust, M.~McDermott, and W.~Lichten.
\newblock Hyperfine structure of the metastable <span class.
\newblock {\em Physical Review}, 120:469, 10 1960.

\bibitem{Faust1961}
W.~L. Faust and M.~N. McDermott.
\newblock Hyperfine structure of the ${(5p)}^{5}(6s)^{3}p_{2}$ state of
  $_{54}\mathrm{Xe}^{129}$ and $_{54}\mathrm{Xe}^{131}$.
\newblock {\em Phys. Rev.}, 123:198--204, Jul 1961.

\bibitem{FERNANDO1960}
P.~C.~B. Fernando, G.~K. Rochester, I.~J. Spalding, and K.~F. Smith.
\newblock The hyperfine structure of 121sb and 123sb.
\newblock {\em https://doi.org/10.1080/14786436008238341}, 5:1291--1298, 1960.

\bibitem{Fisk1997}
Peter~T.H. Fisk, Matthew~J. Sellars, Malcolm~A. Lawn, and Colin Coles.
\newblock Accurate measurement of the 12.6 ghz clock transition in trapped
  171yb+ ions.
\newblock {\em IEEE Transactions on Ultrasonics, Ferroelectrics, and Frequency
  Control}, 44:344--354, 1997.

\bibitem{Frommgen2015}
Nadja Frömmgen, Dimiter~L. Balabanski, Mark~L. Bissell, Jacek Bieroń, Klaus
  Blaum, Bradley Cheal, Kieran Flanagan, Stephan Fritzsche, Christopher
  Geppert, Michael Hammen, Magdalena Kowalska, Kim Kreim, Andreas Krieger,
  Rainer Neugart, Gerda Neyens, Mustafa~M. Rajabali, Wilfried Nörtershäuser,
  Jasna Papuga, and Deyan~T. Yordanov.
\newblock Collinear laser spectroscopy of atomic cadmium.
\newblock {\em The European Physical Journal D 2015 69:6}, 69:1--12, 6 2015.

\bibitem{PerezGalvan2008}
A.~Pérez Galván, Y.~Zhao, and L.~A. Orozco.
\newblock Measurement of the hyperfine splitting of the <span class.
\newblock {\em Physical Review A}, 78:012502, 7 2008.

\bibitem{PerezGalvan2007}
A.~Pérez Galván, Y.~Zhao, L.~A. Orozco, E.~Gómez, A.~D. Lange, F.~Baumer,
  and G.~D. Sprouse.
\newblock Comparison of hyperfine anomalies in the 5 s1 / 2 and 6 s1 / 2 levels
  of 85rb and 87rb.
\newblock {\em Physics Letters, Section B: Nuclear, Elementary Particle and
  High-Energy Physics}, 655:114--118, 11 2007.

\bibitem{Gangrsky1997}
Yu.P. Gangrsky, B.K. Kul’djanov, K.P. Marinova, B.N. Markov, and S.G.
  Zemlyanoi.
\newblock Hyperfine anomaly in the f 3 ds 2 5 l 6 0 , f 3 dsp 7 m 7 and f 3 dsp
  7 l 6 levels in ui.
\newblock {\em Zeitschrift für Physik D Atoms, Molecules and Clusters 1997
  42:1}, 42:1--4, 4 2014.

\bibitem{GOULD1956}
Gordon Gould.
\newblock hfs separations and hfs anomaly in the <span class.
\newblock {\em Physical Review}, 101:1828, 3 1956.

\bibitem{Grossman1999}
J.~S. Grossman, L.~A. Orozco, M.~R. Pearson, J.~E. Simsarian, G.~D. Sprouse,
  and W.~Z. Zhao.
\newblock Hyperfine anomaly measurements in francium isotopes and the radial
  distribution of neutrons.
\newblock {\em Physical Review Letters}, 83:935, 8 1999.

\bibitem{GUSTAVSSON1979}
M.~Gustavsson, G.~Olsson, and A.~Rosén.
\newblock Hyperfine-structure investigation in the 6s5d configuration of135ba
  and137ba.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1979 290:3},
  290:231--243, 9 1979.

\bibitem{Hogervorst1980}
W.~Hogervorst, H.~A. Helms, G.~J. Zaal, J.~Bouma, and J.~Blok.
\newblock The hyperfine structure and nuclear moments of some neutron-deficient
  as isotopes.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1980 294:1},
  294:1--5, 3 1980.

\bibitem{King1956}
J.~H. Holloway, B.~B. Aubrey, and J.~G. King.
\newblock {\em Quarterly Progress Report, Research Laboratory of Electronics,
  Massachusetts Institute of TechnologyApril 15, (unpublished)}, page 35–36,
  1956.

\bibitem{Holloway1962}
W.~W. Holloway, E.~Lüscher, and R.~Novick.
\newblock Hyperfine structure of atomic nitrogen.
\newblock {\em Physical Review}, 126:2109, 6 1962.

\bibitem{Iimura2003}
H.~Iimura, M.~Koizumi, M.~Miyabe, M.~Oba, T.~Shibata, N.~Shinohara, Y.~Ishida,
  T.~Horiguchi, and H.~A. Schuessler.
\newblock Nuclear moments and isotope shifts of <span class.
\newblock {\em Physical Review C}, 68:054328, 11 2003.

\bibitem{Kuhnert1983}
A.~Kuhnert, A.~Nunnemann, and D.~Zimmermann.
\newblock Investigation of the hyperfine structure and isotope shift of the
  542.2 nm line of lu.
\newblock {\em Journal of Physics B: Atomic and Molecular Physics},
  16:4299--4303, 1983.

\bibitem{LURIO1956}
A.~Lurio and A.~G. Prodell.
\newblock Hfs separations and hfs anomalies in the <span class.
\newblock {\em Physical Review}, 101:79, 1 1956.

\bibitem{LUTZ1970}
O.~Lutz.
\newblock The hyperfine structure anomaly of 79br and 81br.
\newblock {\em Physics Letters A}, 31:384, 4 1970.

\bibitem{LUTZ1981}
O.~Lutz, W.~Messner, K.~R. Mohn, and P.~Kroneck.
\newblock Measurement of50v and51v nuclear moments by nuclear magnetic
  resonance methods.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1981 300:2},
  300:111--114, 6 1981.

\bibitem{LUTZ1971}
O.~Lutz, A.~Nolle, and A.~Uhl.
\newblock The hyperfine structure anomalies of67ga,69ga,71ga,72ga,113in,115in,
  and117min.
\newblock {\em Zeitschrift für Physik A Hadrons and nuclei 1971 248:2},
  248:159--161, 3 2005.

\bibitem{LUTZ1978}
O.~Lutz, H.~Oehler, and P.~Kroneck.
\newblock 63cu and65cu fourier transform nuclear magnetic resonance studies.
\newblock {\em Zeitschrift für Physik A Atoms and Nuclei 1978 288:1},
  288:17--21, 3 1978.

\bibitem{moon2009}
Han~Seb Moon, Won-Kyu Lee, and Ho~Suhng Suh.
\newblock Hyperfine-structure-constant determination and absolute-frequency
  measurement of the rb <span class.
\newblock {\em Physical Review A}, 79:062503, 6 2009.

\bibitem{Munch1987}
A.~Münch, M.~Berkler, Ch. Gerz, D.~Wilsdorf, and G.~Werth.
\newblock Precise ground-state hyperfine splitting in <span class.
\newblock {\em Physical Review A}, 35:4147, 5 1987.

\bibitem{Papuga2014}
J.~Papuga, M.~L. Bissell, K.~Kreim, C.~Barbieri, K.~Blaum, M.~De Rydt,
  T.~Duguet, R.~F.~Garcia Ruiz, H.~Heylen, M.~Kowalska, R.~Neugart, G.~Neyens,
  W.~Nörtershäuser, M.~M. Rajabali, R.~Sánchez, N.~Smirnova, V.~Somà, and
  D.~T. Yordanov.
\newblock Shell structure of potassium isotopes deduced from their magnetic
  moments.
\newblock {\em Physical Review C}, 90:034321, 9 2014.

\bibitem{Persson2018Pb}
J~R Persson.
\newblock Hyperfine structure and hyperfine anomaly in pb.
\newblock {\em Journal of Physics Communications}, 2:055028, 5 2018.

\bibitem{Persson2018GdNd}
Jonas~R. Persson.
\newblock Hyperfine anomalies in gd and nd.
\newblock {\em Atoms}, 6:63, 11 2018.

\bibitem{Persson2020}
Jonas~R. Persson.
\newblock Hyperfine anomaly in eu isotopes and the universiability of the
  moskowitz–lombardi formula.
\newblock {\em Atoms}, 8(1), 2020.

\bibitem{Platen1971}
Ch. Platen, J.~Bonn, U.~Köpf, R.~Neugart, and E.W. Otten.
\newblock Spin exchange polarization and hfs anomaly measurement of
  $beta$-active37k.
\newblock {\em Z. Phys.}, 244:44, 2 1971.

\bibitem{Ranjit2014}
G.~Ranjit, D.~Kealhofer, G.~D. Vukasin, and P.~K. Majumder.
\newblock Measurement of <span class.
\newblock {\em Physical Review A}, 89:012511, 1 2014.

\bibitem{Reimann1973}
R.~J. Reimann and M.~N. McDermott.
\newblock Precision magnetic moment determinations for 43-min <span class.
\newblock {\em Physical Review C}, 7:2065, 5 1973.

\bibitem{SCHMELLING1974}
Stephen~G. Schmelling.
\newblock Hyperfine structure in the metastable $d$ states of atomic barium.
\newblock {\em Phys. Rev. A}, 9:1097--1102, Mar 1974.

\bibitem{Schmelling1967}
Stephen~G. Schmelling, Vernon~J. Ehlers, and Howard~A. Shugart.
\newblock Nuclear magnetic moment, hyperfine structure, and hyperfine-structure
  anomaly of <span class.
\newblock {\em Physical Review}, 154:1142, 2 1967.

\bibitem{Schmelling1970}
Stephen~G. Schmelling, Vernon~J. Ehlers, and Howard~A. Shugart.
\newblock Nuclear magnetic moment, hyperfine structure, and hyperfine-structure
  anomaly of <span class.
\newblock {\em Physical Review C}, 2:225, 7 1970.

\bibitem{STINSON1971}
G.~M. Stinson, A.~R. Pierce, J.~C. Waddington, and R.~G. Summers-Gill.
\newblock Spin and magnetic moment of silver-109m.
\newblock {\em https://doi.org/10.1139/p71-109}, 49:906--913, 4 1971.

\bibitem{STROKE1957}
H.~H. Stroke, V.~Jaccarino, Jr~D.~S. Edmonds, and R.~Weiss.
\newblock Magnetic moments and hyperfine-structure anomalies of <span class.
\newblock {\em Physical Review}, 105:590, 1 1957.

\bibitem{THADDEUS1963}
P.~Thaddeus and M.~N. McDermott.
\newblock Level crossings in the <span class.
\newblock {\em Physical Review}, 132:1186, 11 1963.

\bibitem{Trapp2000}
S.~Trapp, G.~Marx, G.~Tommaseo, A.~Klaas, A.~Drakoudis, G.~Revalde,
  G.~Szawiola, and G.~Werth.
\newblock Hyperfine structure and $g$ factor measurements on ba+ and eu+
  isotopes.
\newblock {\em Hyperfine Interactions 2000 127:1}, 127:57--64, 2000.

\bibitem{Wang2014}
Jie Wang, Huifeng Liu, Guang Yang, Baodong Yang, and Junmin Wang.
\newblock Determination of the hyperfine structure constants of the <span
  class.
\newblock {\em Physical Review A}, 90:052505, 11 2014.

\bibitem{Wannberg1970}
B~Wannberg, J~O Jönsson, and L~Sanner.
\newblock Nuclear magnetic moment and hyperfine structure anomaly of 103ag and
  nuclear spin of 13 m 102ag.
\newblock {\em Physica Scripta}, 1:238, 5 1970.

\bibitem{Witte2002}
S.~Witte, E.J. van Duijn, R.~Zinkstok, and W.~Hogervorst.
\newblock High-resolution lif measurements on hyperfine structure and isotope
  shifts in various states of lu i using the second and third harmonic of a cw
  ti:sapphire laser.
\newblock {\em The European Physical Journal D - Atomic, Molecular, Optical and
  Plasma Physics 2002 20:2}, 20:159--164, 8 2002.

\bibitem{WORLEY1965}
Richard~D. Worley, Vernon~J. Ehlers, William~A. Nierenberg, and Howard~A.
  Shugart.
\newblock Hyperfine-structure separation, nuclear magnetic moment, and
  hyperfine-structure anomaly of cesium-131.
\newblock {\em Physical Review}, 140:B1483, 12 1965.

\bibitem{Yordanov2013}
D.~T. Yordanov, D.~L. Balabanski, J.~Bieroń, M.~L. Bissell, K.~Blaum,
  I.~Budinčević, S.~Fritzsche, N.~Frömmgen, G.~Georgiev, Ch. Geppert,
  M.~Hammen, M.~Kowalska, K.~Kreim, A.~Krieger, R.~Neugart, W.~Nörtershäuser,
  J.~Papuga, and S.~Schmidt.
\newblock Spins, electromagnetic moments, and isomers of <span class.
\newblock {\em Physical Review Letters}, 110:192501, 5 2013.

\bibitem{Yordanov2017}
D~T Yordanov, M~L Bissell, K~Blaum, M~De Rydt, Ch~Geppert, J~Krämer, K~Kreim,
  M~Kowalska, A~Krieger, P~Lievens, R~Neugart, G~Neyens, W~Nörtershäuser, L~V
  Rodr{\'{\i}}guez, R~S{\'{a}}nchez, and P~Vingerhoets.
\newblock Spin and magnetic moment of23mg.
\newblock 44(7):075104, may 2017.

\bibitem{Zhang2015}
J.~Zhang, M.~Tandecki, R.~Collister, S.~Aubin, J.A. Behr, E.Gomez, G.~Gwinner,
  L.A: Orozco, FrPNC collaboration, M.R. Pearson, and G.D. Sprouse.
\newblock Hyperfine anomalies in fr: Boundaries of the spherical single
  particle model.
\newblock {\em Physical Review Letters}, 115:042501, 7 2015.

\end{thebibliography}


\begin{thebibliography}{999}
                                 
\bibitem {otten}E.W. Otten, Treatise on Heavy-Ion Science\textit{\ }vol 8, ed.
D. Allan Bromely (Plenum, New York 1989) 515

\bibitem {billowes}J.Billowes and P. Campbell, J. Phys. G
\textbf{21}, (1995) 707.

\bibitem {Buttgenbach}S. B\"{u}ttgenbach, Hyperfine Int. \textbf{20}, (1984) 1

\bibitem {stroke}H.H. Stroke, V. Jaccarino, D.S. Edmonds, Jr and
R. Weiss, Phys. Rev. \textbf{105}, (1957) 590.

\bibitem {mosko}P.A. Moskowitz, C.H. Liu, G. Fulop and H.H. Stroke, Phys. Rev. \textbf{C4}, (1971) 620

\bibitem{PERSSON201362} Persson, J.R., Atomic Data and Nuclear Data Tables \textbf{99}, (2013) 62-68

\bibitem {Kopf}  H. Kopfermann, Kernmomente (Akademische Verlagsgesellschaft, Leipzig, 1940)

\bibitem {Bitter} F. Bitter, Phys. Rev. \textbf{76},150 (1949)

\bibitem {bohrweisskopf}A. Bohr and V.F. Weisskopf, Phys. Rev. \textbf{77},
(1950) 94

\bibitem {Breit}E. Rosenthal and G. Breit, Phys. Rev. \textbf{41},
(1932) 459

\bibitem {Crawford}M. Crawford and A. Schawlow, Phys. Rev.
\textbf{76}, (1949) 1310

\bibitem {Pallas}N.J. Ionesco-Pallas, Phys. Rev. \textbf{117}, (1960) 505

\bibitem {Rosenberg}H.J. Rosenberg and H.H. Stroke, Phys. Rev. A
\textbf{5}, (1972) 1992

\bibitem{Heggset2020} Heggset, Tarjei and Persson, Jonas R.,Atoms \textbf{8}, (2020) 86



\bibitem{Galvan}
Perez Galvan, A., Zhao,  Y., Orozco, L.A., Gomez E., Lange, A.D., Baumer,A. and
  Sprouse, G.D.: Phys. Lett. \textbf{655B}, 114 (2007)

\bibitem {Ref4}I. Lindgren and J. Morrison, Atomic Many-Body Theory
(Springer-Verlag, Berlin 1983)

\bibitem {Ref5}P.G.H. Sandars and J. Beck, Proc. R. Soc. London \textbf{A289},
97 (1965)

\bibitem {persson}J.R. Persson, Euro. Phys. J. \textbf{A2},(1998) 3

\bibitem{Sanamyan} G. Sanamyan, and B. M. Roberts and J. S. M. Ginges, Phys. Rev. Lett. \textbf{130}, 053001,(2023)

\bibitem{Jonsson} Jönsson, P.; Godefroid, M.; Gaigalas, G.; Ekman, J.; Grumer, J.; Li, W.; Li , J.; Brage, T.; Grant, I.P.; Bieroń, J.; Fischer, C.F. An Introduction to Relativistic Theory as Implemented in GRASP. Atoms 2023, 11, 7. https://doi.org/10.3390/atoms11010007

\bibitem{Papoulia}Papoulia, A. and Carlsson, B. G. and Ekman, J.,Phys. Rev. A,
  \textbf{94},042502 (2016)


\bibitem {stone}N.J. Stone At. Data Nucl. Data Tables \textbf{90} (2005) 75,176



\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{longtable}
\usepackage{lscape}

\usepackage{amsmath}

\usepackage{amssymb}

\begin{document}




\title{{Table of hyperfine anomaly in atomic systems - 2023}}

  \author{J.R. Persson\\
  Department of Physics\\
  NTNU\\
  NO-7491 Trondheim\\
  Norway\\
  E-mail: jonas.persson@ntnu.no}



\date{24.04.2023} 
\maketitle

\begin{abstract}
 This table is an updated compilation of experimental values of the magnetic hyperfine anomaly in atomic and ionic systems. The literature search covers the period up to December 2022. A short discussion on general trends of the hyperfine anomaly and the theoretical developments are performed.
\end{abstract}





\newpage



\section{Introduction}
The effect on the hyperfine structure by nuclear charge and magnetisation distributions has been studied since the first experimental observation in Rb was made in 1949 by Bitter \cite{Bitter}. The effect of an extended nuclear magnetisation is manifested by the difference between the assumed point-like and actual magnetic dipole hyperfine interaction, as first anticipated by Kopfermann \cite{Kopf} towards the end of the 1930ties. The effect of extended magnetisation was calculated by Bohr and Weisskopf \cite{bohrweisskopf} in 1950 and is therefore known as the Bohr-Weisskopf (BW) effect. The extended nuclear charge distribution will, in addition to the Isotope Shift (IS), also give rise to an effect on the hyperfine structure. This is due to a modification of the electronic wavefunctions by the extended nuclear charge distribution, the so-called "Breit-Rosenthal-Crawford-Schawlow", or Breit-Rosenthal(BR) correction \cite{Breit,Crawford,Pallas,Rosenberg}.

These effects are normally quite small in atomic systems and remain quite difficult to calculate accurately, except in one-electron systems. This is the reason why it is more sensible to to study differential differences between isotopes, the so-called differential hyperfine anomaly. Due to the small size of the differential effect only a few systematic investigations have been performed \cite{stroke,mosko,Buttgenbach, PERSSON201362}. This review is an update of the 2013 paper as there have been some developments since it was published.

\section{Hyperfine Anomaly} 

The only way to probe the structural properties of the nucleus is with the use of "penetrating" electrons, that is the electrons wavefunction must have a nonzero probability to be found inside the nucleus. This means that only $s_{1/2}$ and relativistic $p_{1/2}$-electrons can be used. One should note that electron-electron interaction can cause an s- or p- electron-like behaviour, thus giving rise to an apparent nonzero probability for other electrons. A more pronounced effect can be obtained with the use of muons, this is outside the scope of this review.

The magnetic dipole hyperfine interaction is represented by the Hamiltonian
	\begin{equation}
    H=aI \cdot J,
    \end{equation}
Where $a$ is the magnetic dipole hyperfine interaction constant. I and J are the nuclear and electron angular momenta, respectively. In the case of an extended nucleus, the point-like hyperfine interaction constant $a_{point}$, will be modified by two effects:\\
1.	The modification of the electronic wavefunctions by the extended nuclear charge distribution, the "Breit-Rosenthal-Crawford-Schawlow" correction ($\epsilon_{BR}$) \cite{Breit,Crawford,Pallas,Rosenberg}.\\
2.	The extended and distributed nuclear magnetisation, the Bohr-Weisskopf effect,($\epsilon_{BW}$ )\cite{bohrweisskopf}.\\
This yields
    \begin{equation}
    a=a_{point}\left(  1+\epsilon_{BW}\right)\left(  1+\epsilon_{BR}\right)
    \end{equation}
Where $a$ denotes the experimental value of the magnetic dipole hyperfine interaction constant.
The hypothetical $a_{point}$ can normally not be calculated with sufficient precision for ordinary atoms, as is the case in muonic atoms and hydrogen-like ions. However, these uncertainties in point-like interactions cancel if we take the ratio of the $a$ values for two isotopes. That is, we can express the ratio without the need for precise calculations:
	\begin{equation}
    \frac{a_{1}}{a_{2}}=
    {\frac{g_I{(1)}}{g_I{(2)}}
    {\frac{[1+\epsilon_{BW}(1)][1+\epsilon_{BR}(1)]}{[1+\epsilon_{BW}(2)][1+\epsilon_{BR}(2)]}}}
    \end{equation}

Where ${g_I = -\mu_I/I}$ is the nuclear gyromagnetic ratio. As we can assume, in the case of electrons, that $\epsilon$  are generally small, we can rewrite (3) as:
    \begin{eqnarray}
    \frac{a_{1}}{a_{2}}\approx{\frac{g_I{(1)}}{g_I{(2)}}{[1+\epsilon_{BW}(1)-\epsilon_{BW}(2)][1+\epsilon_{BR}(1)-\epsilon_{BR}(2)]}}\nonumber\\ ={\frac{g_I{(1)}}{g_I{(2)}}{[1+^{1}\Delta^{2}_{BW}][1+^{1}\Delta^{2}_{BR}]}}
    \end{eqnarray}
	
Here we define the differential hyperfine anomaly of Bohr-Weisskopf and Breit-Rosenthal corrections, respectively as:
    \begin{equation}
    ^{1}\Delta^{2}_{BW}\equiv \epsilon_{BW}(1)-\epsilon_{BW}(2)
    \end{equation}
    \begin{equation}
    ^{1}\Delta^{2}_{BR}\equiv \epsilon_{BR}(1)-\epsilon_{BR}(2)
    \end{equation}
	
Calculations of $\epsilon_{BR} $ may show significant values, while the differential $^{1}\Delta^{2}_{BR}$ is normally expected to be small and negligible compared to $^{1}\Delta^{2}_{BW}$\cite{Rosenberg}. However, in cases where the nuclei are similar with respect to nuclear spin and nuclear configuration $^{1}\Delta^{2}_{BR}$ can be on the same order as $^{1}\Delta^{2}_{BW}$ \cite{Heggset2020} and should be calculated. 

Since the BR effect can be neglected or calculated the subscripts can be dropped obtaining
    \begin{equation}
    \frac{a_{1}}{a_{2}}\approx{\frac{g_I{(1)}}{g_I{(2)}}{[1+^{1}\Delta^{2}]}}
    \end{equation}
	
In most cases will the hyperfine anomaly be of the order of $10^{-3}$ , making it  necessary to know the hyperfine interaction constants, $a$, and the nuclear gyromagnetic values ($\frac{\mu_I}{I}$) with at least an accuracy of $10^{-4}$ or better to obtain values accurate to 10\% for the hyperfine anomaly\cite{Buttgenbach}. Precision values of the hyperfine interaction constants, $a$, and independently measured nuclear magnetic dipole moments ($\mu_I)$ are thus needed to obtain the differential hyperfine anomaly,$^{1}\Delta^{2}$.

\section{State-dependent and s-electron hyperfine anomaly}
It has been observed that the hyperfine anomaly is depending on the atomic state, where the values for different states can vary significantly, but generally shows an n-independence, as found in Rb \cite{Galvan}. It should be noted that the n-independence is the case in single s- and p-electron states.  While the hyperfine anomalies normally are on the order of 1\% or less, the state-dependent hyperfine anomaly can attain values up to 10 \% in special cases. The origin of this can be found in the hyperfine interaction operators and different contributions to the hyperfine interaction.

The hyperfine interaction can be represented by the following operators\cite{Ref4,Ref5}:
\begin{equation}
h = \frac{{\mu _0 }}{{4\pi }}2\mu _B \sum\limits_{i = 1}^N {\left[ {{\mathop{\rm l}\nolimits} \left\langle {r^{ - 3} } \right\rangle ^{01}  - \sqrt {10} \left( {{\mathop{\rm sC^2}\nolimits}  } \right)^1 \left\langle {r^{ - 3} } \right\rangle ^{12}  + {\mathop{\rm s}\nolimits} \left\langle {r^{ - 3} } \right\rangle ^{10} } \right]} _i  \cdot \mu _I,
\end{equation}
where ${{\mathop{\rm l}\nolimits}}$ and ${\mathop{\rm s}\nolimits}$ are the orbital and spin angular momentum operators, respectively, of the electron, ${{\mathop{\rm sC^2}\nolimits}  }$ is a tensor product of ${\mathop{\rm s}\nolimits}$ and ${\mathop{\rm C^2}\nolimits}$ of rank 1. The indices stand for the rank in the spin and orbital spaces, respectively. The hyperfine interaction can be divided into three parts, orbital, spin-dipole and contact (spin) interaction, where only the contact (spin) interaction contributes to the hyperfine anomaly. This means that only s and $p_{1/2}$ electrons contribute to the hyperfine anomaly through the contact (spin) interaction. We can thus write the $a$ constant as consisting of a contact and non-contact part:
\begin{equation}
a=a_{nc}+a_{c},
\end{equation}

Where $a_{c}$ is the contribution due to the contact interaction and $a_{nc}$ the contribution due to non-contact interactions. The experimental hyperfine anomaly, defined with the experimental $a$ constant, should then be rewritten to obtain the contact contribution to the hyperfine anomaly:
\begin{equation}
{^{1}\Delta_{exp}^{2}}={^{1}\Delta_{c}^{2}}{\frac{a_{c}}{a}}
\end{equation}
where $^{1}\Delta^{2}_{c}$ is the hyperfine anomaly due to the contact interaction, that is, for an $s$- or $p_{1/2}$-electron. The hyperfine anomaly is most often given as the experimental state-dependent hyperfine anomaly, as the contact (s-electron) anomaly can be difficult to extract in the case of many-electron configurations.

However, observing that the contact anomaly $^{1}\Delta_{c}^{2}$ is a property that is independent of the state and constant within the atom. One finds that it is possible to extract the anomaly solely from the $a$-constants of two different atomic states \cite{persson}, provided the ratio $a_c/a$ differs substantially for the states. Comparing the ratio of $a$-constants for two isotopes in two atomic levels then gives:
\begin{equation}
{\frac{{a_{B}^{(1)}/a_{B}^{(2)}}}{{a_{C}^{(1)}/a_{C}^{(2)}}}}\approx
{1+{^{1}\Delta_{c}^{2}}({{\frac{a_{c}^{B}}{a^{B}}}-{\frac{a_{c}^{C}}{{a^{C}}}
})}}
\end{equation}
Where B and C denote different atomic levels and 1 and 2 denote different isotopes. The ratio between the two $a$-constant ratios ($\frac{a_{c}^{X}}{a^{X}}$) for the isotopes will only depend on the  difference in the contact contributions of the two atomic levels and the hyperfine anomaly for the s-electron and will be the same for all isotopes. It should be pointed out that the atomic states used must differ significantly in the ratio $\frac{a_{c}}{a}$, as a small difference will lead to an increased sensitivity to errors \cite{persson}. 
If we have a state where the hyperfine anomaly is expected to be small, the ${\frac{a_{c}^{C}}{{a^{C}}}}$ can be set to zero for that state, thus serving as a substitute for $g_I$. It is then possible to obtain values of the contact hyperfine anomaly for level B. 
Since the contact interaction is proportional to the $g_J$-factor of the state, it is easy to identify suitable candidates from the values of the theoretical or experimental $g_J$-factor . The optimal case would be two atomic levels within the same multiplet where the experimental (or theoretical) $g_J$ is greater and smaller than 1, respectively.
This method is especially useful for unstable isotopes where high-precision measurements of the nuclear magnetic moment do not exist. 
Furthermore, states with a substantial difference in  the ratio $a_c/a$ are also preferable for studies of the isotope shift. If measurements are performed on more than three atomic levels, it is also possible to deduce the nuclear magnetic moment ratio and the hyperfine anomaly simultaneously. This will give the nuclear magnetic moment with high accuracy provided the nuclear magnetic moment is known for at least one stable isotope.
The ratio $a_c/a$ can also be calculated with high accuracy making this method very useful where calculations can be done. 

\section{Calculation related to hyperfine anomaly}
Calculations of the BW-effect and subsequently hfa have long been performed by modelling the nuclear magnetisation using a uniform or Fermi distribution of magnetisation. This combined with difficulties with the accuracy of wavefunctions close and within the nucleus for many-electron systems, has given values far from experimental values. However, progress, both theoretically and experimentally, has made it possible to link few-electron (highly charged ions) and muonic systems to calculations of many-electron systems \cite{Sanamyan}. This development has been a by-product of the effort for finding new physics in precision atomic searches \cite{Sanamyan}.  
The work of Sanamyan et al. \cite{Sanamyan} also showed that the simplistic model of uniform magnetisation do not describe the BW-effect in a good way and should be replaced with other available models such as different  varieties of the single-particle model.
Simultaneously, work on improving the GRASP code is on the way \cite{Jonsson} including the possibility to calculate the BW-effect. This will open possibilities for a systematic studies as it will become easier to compare experimental values with own calculations. In addition, as was found by Heggset et. al\cite{Heggset2020} the BR-effect can be calculated readily using the present GRASP code, it is still interesting to study the effect using a more realistic charge distribution for heavy nuclei \cite{Papoulia}.

\section{Trends in hyperfine anomaly and need for more measurements}
Even if the hfa has been studied since the 1950ties, the values for different elements are sporadic and very few systematic studies have been done due to the requirements of high-precision measurements of both hyperfine structure and nuclear magnetic dipole moments. It is well established that the hfa (BW-effect) attains rather similar values within the same nuclear configuration and spin as has been observed in Fr (I=9/2), Hg (I=13/2), Pb (I=13/2) and Eu (I=5/2). In these cases, the BR-effect may be on the same order as the BW-effect \cite{Heggset2020} and can not be neglected. Since it is possible to calculate the BR-effect it is possible to assume a smaller uncertainty in the determination of the nuclear magnetic dipole moment of unstable isotopes if the hfa (BW- and BR-effect) is known in an isotope with the same nuclear configuration. This is of special importance in laser spectroscopy studies in elements with $Z<83$.
The procedure to use the Moscowitz-Lombardi semi-empirical formula as a tool to estimate the hfa is questionable, as it has been shown that it is not universal and only seems to be valid in the Z=80 region. However this should be studied further with more systematic measurements in this region.
There is also a need for more systematic studies in the vicinity of the alkali elements as these are used in precision atomic calculations with the aim to find new physics \cite{Sanamyan}.

\section{Outlook}
Even if the hyperfine anomaly and BW-effect has been a small subject, progress both experimentally and in calculations open the possibility for more systematic studies with better understanding of nuclear structure as well as a means to test the atomic wavefunctions in precision atomic calculations. From that perspective one can expect a renewed interest and more precision measurements of the hyperfine structure and from these, the hyperfine anomaly.

\section{Policies followed in the compilation}

The hyperfine anomaly is given with the lightest stable isotope as the reference isotope. The lightest naturally abundant isotope was used for U and the designation of the original article was used for Fr. In most cases the original article, where the hyperfine anomaly has been derived, is used. In the case of recent, more precise values of the nuclear magnetic moment, the hyperfine anomaly has been updated, accordingly. The nuclear magnetic moments of Stone \cite{stone} have been used, unless more precise values of ratios are available. Special care was taken to use magnetic moments obtained by the same method.
The hyperfine anomaly is given as state-dependent if not stated otherwise. If the s-electron hyperfine anomaly is known, no extensive listing of state-dependent hyperfine anomaly is given, unless these are of special interest.




\clearpage

\section*{Table 1.\label{tbl1te} Experimental data of hyperfine anomaly values in atomic systems}

\begin{center}
\begin{tabular}{ll}

Element & The element studied\\

Isotope 1
	& Reference isotope for the hyperfine anomaly\\
Spin 1
	& Spin for isotope 1\\

Isotope 2
	& The second isotope used.\\

Spin 2
	& Spin for isotope 2\\

State-/s- anomaly	& The atomic state for which the experimental hfa\\
                        & has been determined or the s-electron hfa.\\

$^{1}\Delta^{2}(\%)$
	& Hyperfine anomaly given in \%.\\

Reference
	& Original article where $^{1}\Delta^{2}(\%)$ or the experimental hyperfine\\
    & interactions constants is given.\\

\end{tabular}
\end{center}
\label{tableI}


\bigskip

\newpage

\begin{landscape}




\setlength{\LTleft}{0pt}
\setlength{\LTright}{0pt}


\setlength{\tabcolsep}{0.5\tabcolsep}

\renewcommand{\arraystretch}{1.0}

\footnotesize 

\begin{longtable}{lcccclcr}
\caption{Experimental data of hyperfine anomaly values in atomic systems}\\
\mbox{Element} & \mbox{isotope 1} & \mbox{spin 1}& \mbox{isotope 2} & \mbox{spin 2}& Atomic state/ s-anomaly & $^{1}\Delta^{2}(\%)$ & Reference\\
\hline\\
\endfirsthead\\
\caption[]{(continued)}\\
\mbox{Element} & \mbox{isotope 1} & \mbox{spin 1}& \mbox{isotope 2} & \mbox{spin 2}& Atomic state/ s-anomaly & $^{1}\Delta^{2}(\%)$ & Reference\\
\hline\\
\endhead

Li & 6 & 1 & 7 & 3/2 & 2s $^{2}$S$_{1/2}$, s-anomaly & 0.0068067(8) & \cite{Allegrini22} \\ 
\\
&  &&&  & 3s $^{2}$S$_{1/2}$, s-anomaly & 0.054(21) & \cite{Allegrini22} \\
&  &&&  & 2p $^{2}$P$_{1/2}$ & -0.1734(2) & \cite{Allegrini22} \\
&  &&&  & 2p $^{2}$P$_{3/2}$ & -0.155(8) & \cite{Allegrini22} \\
\\

N & 14 & 1 & 15 & 1/2 & 2p$^{3}$ $^{4}$S$_{3/2}$ & 0.0999(4) & \cite{Holloway1962} \\
\\

Na & 23 & 3/2 & 24 & 4 & 3s $^{2}$S$_{1/2}$, s-anomaly & 0.0013(30) & \cite{Beckmann1974}\cite{Chan1966} \\
\\

Mg & 25 & 5/2 & 23 & 3/2 & 3s $^{2}$S$_{1/2}$ - 3p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & -0.04(8) & \cite{Yordanov2017}\\
\\

Cl & 35 & 3/2 & 37 & 3/2 & 3p$^{5}$ $^{2}$P$_{3/2}$ & -0.00381(2) & \cite{King1956}
\\
\\

K & 39 & 3/2 & 37 & 3/2 & 4s $^{2}$S$_{1/2}$, s-anomaly & -0.249(35) & \cite{Platen1971}
\\
& 39 & 3/2 & 40 & 4 & 4s $^{2}$S$_{1/2}$, s-anomaly & 0.467(2) & \cite{Allegrini22}
\\

& 39 & 3/2 & 41 & 3/2 & 4s $^{2}$S$_{1/2}$, s-anomaly & -0.22937(13) & \cite{Beckmann1974}\cite{Allegrini22} \\

& 39 & 3/2 & 42 & 2 & 4s $^{2}$S$_{1/2}$, s-anomaly & 0.336(38) & \cite{Chan1969} \\

 & 39 & 3/2 & 38 & 3 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.53(44) & \cite{Papuga2014}\\
 & 39 & 3/2 & 40 & 4 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.43(17) & \cite{Papuga2014}\\
 & 39 & 3/2 & 41 & 3/2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & -0.23(31) & \cite{Papuga2014}\\
 & 39 & 3/2 & 42 & 2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.99(36) & \cite{Papuga2014}\\
 & 39 & 3/2 & 44 & 2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.47(47) & \cite{Papuga2014}\\
 & 39 & 3/2 & 46 & 2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.40(39) & \cite{Papuga2014}\\
 & 39 & 3/2 & 47 & 1/2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta_{p}$) & 0.28(16) & \cite{Papuga2014}\\
 & 39 & 3/2 & 48  & 1 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.57(35) & \cite{Papuga2014}\\
 & 39 & 3/2 & 49 & 1/2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.24(29) & \cite{Papuga2014}\\
 & 39 & 3/2 & 51 & 3/2 &  4s $^{2}$S$_{1/2}$ - 4p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.57(250) & \cite{Papuga2014}\\

\\
V & 50 & 6 & 51 & 7/2 & 3d$^{3}$4s$^{2}$ $^{4}$F$_{5/2}$ & 0.0007(10) & \cite{LUTZ1981}
\\
&  &&&  & 3d$^{4}$4s $^{6}$D$_{1/2}$ & 0.034(60) & \cite{Cochrane1998} \\
\\

Cu & 63 & 3/2 & 65 & 3/2 & 3d$^{10}$4s $^{2}$S$_{1/2}$ & 0.004861(9) & \cite{LUTZ1978}
\\
&  &&&  & 3d$^{9}$4s4p $^{4}$P$_{5/2}$ & 0.00340(11) & \cite{LUTZ1978}\cite%
{BLACHMAN1969} \\
&  &&&  & 3d$^{9}$4s4p $^{4}$P$_{9/2}$ & 0.00305(17) & \cite{LUTZ1978}\cite%
{BLACHMAN1969} \\
\\

Ga & 69 & 3/2 & 67 & 3/2 & 4p $^{2}$P$_{1/2}$ & -0.00050(7) & \cite{LUTZ1971} \\
&  & & &  & 4p $^{2}$P$_{3/2}$ & 0.00200(16) & \cite{LUTZ1971} \\
& 69 & 3/2 & 71 & 3/2 & 4p $^{2}$P$_{1/2}$ & 0.00063(6) & \cite{LUTZ1971} \\
&  & & &  & 4p $^{2}$P$_{3/2}$ & -0.00252(12) & \cite{LUTZ1971} \\
& 71 & 3/2 & 72 & 3 & 4p $^{2}$P$_{1/2}$ & 0.0043(6) & \cite{LUTZ1971} \\
&  & & &   & 4p $^{2}$P$_{3/2}$ & -0.0170(18) & \cite{LUTZ1971} \\
\\

As & 75 & 3/2 & 70 & 4 & 4p$^{3}$ $^{4}$S$_{3/2}$ & -0.35(2) & \cite{Hogervorst1980} \\
\\
Br & 79 & 3/2 & 81 & 3/2 & 4p$^{5}$ $^{2}$P$_{3/2}$ & -0.00003(4) & \cite{BROWN1966}\cite%
{LUTZ1970} \\
\\
Rb & 85 & 5/2 & 84 & 2 & 5s $^{2}$S$_{1/2}$, s-anomaly & -1.7(1.0) & \cite%
{ACKERMANN1973} \\
& 85 & 5/2 & 86 & 2 & 5s $^{2}$S$_{1/2}$, s-anomaly & 0.17(9) & \cite{BRASLAU1961} \\
& 85 & 5/2 & 87 & 3/2 & 5s $^{2}$S$_{1/2}$, s-anomaly & 0.35141(2) & 
\cite{Allegrini22} \\
&  &&&  & 6s $^{2}$S$_{1/2}$, s-anomaly & 0.361(19) & \cite{Allegrini22}\cite{PerezGalvan2007}\cite{PerezGalvan2008} \\
&  &&&  & 7s $^{2}$S$_{1/2}$, s-anomaly & 0.342(3) & \cite{Allegrini22} \\
&  &&&  & 5p $^{2}$P$_{1/2}$ & 0.55(8) & \cite{Allegrini22}\\
&  &&&  & 5p $^{2}$P$_{3/2}$ & 0.168(5) & \cite{Allegrini22} \\
&  &&&  & 6p $^{2}$P$_{1/2}$ & 0.31(7) & \cite{Allegrini22} \\
&  &&&  & 6p $^{2}$P$_{3/2}$ & 0.46(5) & \cite{Allegrini22} \\

&  &&&  & 4d $^{2}$D$_{3/2}$ & 0.347(4) & \cite{moon2009} \\
&  &&&  & 4d $^{2}$D$_{5/2}$ & 0.41(9) & \cite{Wang2014} \\
&  &&&  & 4d $^{2}$D$_{5/2}$ & 0.60(15) & \cite{Allegrini22} \\
&  &&&  & 5d $^{2}$D$_{3/2}$ & 0.279(6) & \cite{Allegrini22} \\
&  &&&  & 5d $^{2}$D$_{5/2}$ & 0.44(5) &\cite{Allegrini22} \\

\\
Mo & 95 & 5/2 & 97 & 5/2 & 4d$^{5}$5s $^{7}$S$_{3}$ & -0.0101(2) & \cite{Buttgenbach1974}
\\
\\
Ru & 99 & 5/2 & 101 & 5/2 & s-anomaly & -0.0173(1) & \cite{Buttgenbach1977} \\
\\

Ag & 107 & 1/2 & 103 & 7/2 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -3.4(1.7) & \cite{Wannberg1970} \\
& 107 & 1/2 & 108 & 1 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -2.6(7) & \cite{CUSSENS1969} \\
& 107 & 1/2 & 109 & 1/2 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -0.41274(29) & \cite{DAHMEN1967}
\\
& 107 & 1/2 & 109$^{m}$ & 7/2 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -3.8(4.1) & \cite%
{STINSON1971} \\
&  &&&  &  & -0.85(1.19) & \cite{STINSON1971}, $\mu _{I}$ from \cite{EDER1985}
\\
& 107 & 1/2 & 110 & 1 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -3.1(1.4) & \cite{CUSSENS1969}
\\
& 107 & 1/2 & 110$^{m}$ & 6 & 4d$^{10}$5s $^{2}$S$_{1/2}$ & -2.88(13) & \cite{Schmelling1967}\\
\\

Cd & 111 & 1/2 & 107 & 1/2 & 5s $^{2}$S$_{1/2}$ & -0.14(3) & \cite{Yordanov2013} \\
& 111 & 1/2 & 109 & 5/2& 5s $^{2}$S$_{1/2}$ & -0.14(3) & \cite{Yordanov2013} \\
& 111 & 1/2 & 113 & 1/2 & 5s $^{2}$S$_{1/2}$ & -0.02(3) & \cite{Yordanov2013} \\
& 111 & 1/2 & 113$^{m}$ & 11/2& 5s $^{2}$S$_{1/2}$ & -0.13(3) & \cite{Yordanov2013} \\
& 111 & 1/2 & 115 & 1/2 & 5s $^{2}$S$_{1/2}$ & 0.02(3) & \cite{Yordanov2013} \\
& 111 & 1/2 & 115$^{m}$ & 11/2& 5s $^{2}$S$_{1/2}$ & -0.08(3) & \cite{Yordanov2013} \\
\\

& 111 & 1/2 & 107 & 1/2 & 5s6s $^{3}$S$_{1}$ & -0.18(4) & \cite{Frommgen2015} \\
& 111 & 1/2 & 109 & 5/2 & 5s6s $^{3}$S$_{1}$ & -0.12(1) & \cite{Frommgen2015} \\
& 111 & 1/2 & 111$^{m}$ & 11/2 & 5s6s $^{3}$S$_{1}$ & -0.10(4) & \cite{Frommgen2015} \\
& 111 & 1/2 & 113 & 1/2 & 5s6s $^{3}$S$_{1}$ & -0.01(1) & \cite{Frommgen2015} \\
& 111 & 1/2 & 113$^{m}$ & 11/2& 5s6s $^{3}$S$_{1}$ & -0.08(1) & \cite{Frommgen2015} \\
& 111 & 1/2 & 115 & 1/2 & 5s6s $^{3}$S$_{1}$ & 0.05(3) & \cite{Frommgen2015} \\
& 111 & 1/2 & 113$^{m}$ & 11/2& 5s6s $^{3}$S$_{1}$ & -0.09(2) & \cite{Frommgen2015} \\
\\
& 111 & 1/2 & 107 & 1/2 & 5s5p $^{3}$P$_{1}$ & -0.0958(8) & \cite{THADDEUS1963} \\
& 111 & 1/2 & 109 & 5/2 & 5s5p $^{3}$P$_{1}$ & -0.0912(7) & \cite{THADDEUS1963} \\
& 111 & 1/2 & 113 & 1/2 & 5s5p $^{3}$P$_{1}$ & -0.00023(40) & \cite{CHANEY1969} \\
& 111 & 1/2 & 113$^{m}$ & 11/2& 5s5p $^{3}$P$_{1}$ & -0.0773(5) & \cite{CHANEY1969} \\
& 111 & 1/2 & 115 & 1/2 & 5s5p $^{3}$P$_{1}$ & 0.244(65) & \cite{CHANEY1969} \\
& 111 & 1/2 & 115$^{m}$ & 11/2& 5s5p $^{3}$P$_{1}$ & -0.236(90) & \cite{CHANEY1969} \\
\\
& 111 & 1/2 & 107 & 1/2 & 5s5p $^{3}$P$_{2}$ & -0.17(5) & \cite{Frommgen2015} \\
& 111 & 1/2 & 109 & 5/2 & 5s5p $^{3}$P$_{2}$ & -0.12(1) & \cite{Frommgen2015} \\
& 111 & 1/2 & 111$^{m}$ & 11/2 & 5s5p $^{3}$P$_{2}$ & -0.08(4) & \cite{Frommgen2015} \\
& 111 & 1/2 & 113 & 1/2 & 5s5p $^{3}$P$_{2}$ & -0.00143(6) & \cite{FAUST1960} \\
& 111 & 1/2 & 113 & 1/2 & 5s5p $^{3}$P$_{2}$ & 0.00(1) & \cite{Frommgen2015} \\
& 111 & 1/2 & 113$^{m}$ & 11/2& 5s5p $^{3}$P$_{2}$ & -0.08(2) & \cite{Frommgen2015} \\
& 111 & 1/2 & 115 & 1/2 & 5s5p $^{3}$P$_{2}$ & 0.01(2) & \cite{Frommgen2015} \\
& 111 & 1/2 & 115$^{m}$ & 11/2 & 5s5p $^{3}$P$_{2}$ & -0.09(4) & \cite{Frommgen2015} \\
\\

In & 113 & 9/2 & 115 & 9/2 & 5p $^{2}$P$_{1/2}$ & 0.00075(13) & \cite{ECK1957} \\
&  &&&  & 5p $^{2}$P$_{3/2}$ & -0.00238(13) & \cite{ECK1957} \\
\\
Sn & 115 & 1/2 & 117 & 1/2 & 5p$^{2}$ $^{3}$P$_{1}$ & 0,0034(10) & \cite{CHILDS1965} \\
&  &&&  & 5p$^{2}$ $^{3}$P$_{2}$ & -0.0003(10) & \cite{CHILDS1965} \\
& 117 & 1/2 & 119 & 1/2 & 5p$^{2}$ $^{3}$P$_{1}$ & 0.0049(10) & \cite{CHILDS1965} \\
&  &&&  & 5p$^{2}$ $^{3}$P$_{2}$ & -0.0009(10) & \cite{CHILDS1965} \\
&  &&&  & 5p$^{2}$ $^{1}$D$_{2}$ & +0.0001(10) & \cite{CHILDS1971} \\
\\
Sb & 121 & 5/2 & 123 & 7/2 & 5p$^{3}$ $^{4}$S$_{3/2}$ & -0.323(9) & \cite{FERNANDO1960}
\\
\\

Xe & 129 & 1/2 & 131 & 3/2 & 6s $^{2}$S$_{1/2}$, s-anomaly & 0.0440(44) & \cite{Faust1961}
\\
\\

Cs & 133 & 7/2 & 131 & 5/2 & 5p$^{5}$6s $^{3}$P$_{2}$, s-anomaly & 0.45(5) & \cite{WORLEY1965}
\\
& 133 & 7/2 & 134 & 4 & 6s $^{2}$S$_{1/2}$, s-anomaly & 0.169(30) & \cite{STROKE1957} \\
& 133 & 7/2 & 134$^{m}$ & 8 & 6s $^{2}$S$_{1/2}$, s-anomaly & -1.38(3) & \cite%
{COHEN1962} \\
& 133 & 7/2 & 135 & 7/2 & 6s $^{2}$S$_{1/2}$, s-anomaly & 0.037(9) & \cite{STROKE1957} \\
& 133 & 7/2 & 137 & 7/2 & 6s $^{2}$S$_{1/2}$, s-anomaly & 0.0018(40) & \cite{STROKE1957}
\\
\\
Ba & 135 & 3/2 & 137 & 3/2 & 5d6s $^{3}$D$_{1}$ & -0.205(7) & \cite{GUSTAVSSON1979} \\
&  &&&  & 5d6s $^{3}$D$_{2}$ & -0.179(22) & \cite{GUSTAVSSON1979} \\
&  &&&  & 5d6s $^{3}$D$_{3}$ & -0.188(17) & \cite{GUSTAVSSON1979} \\
&  &&&  & 5d6s $^{1}$D$_{2}$ & -0.212(26) & \cite{SCHMELLING1974} \\
&  &&&  & Ba$^{+}$ 6s $^{2}$S$_{1/2}$, s-anomaly & -0.191(5) & \cite{Trapp2000}
\\
\\
La & 138 & 5 & 139 & 7/2 & 5d6p $^{3}$D$_{1}$ & -0.35(23) & \cite{Iimura2003}\cite%
{CHILDS1979} \\
\\

Nd & 143 & 7/2 & 145 & 7/2 & s-anomaly & 0.2034(63) & \cite{Persson2018GdNd} \\
\\
Eu & 151 & 5/2 & 145 & 5/2 & s-anomaly & -0.08(15) & \cite{Persson2020} \\
& 151 & 5/2 & 146 & 4 & s-anomaly & 0.12(50) & \cite{Persson2020} \\
& 151 & 5/2 & 147 & 5/2 & s-anomaly & -0.12(17) & \cite{Persson2020} \\
& 151 & 5/2 & 148 & 5 & s-anomaly & 0.08(31) & \cite{Persson2020} \\
& 151 & 5/2 & 149 & 5/2 & s-anomaly & -0.19(16) & \cite{Persson2020} \\
& 151 & 5/2 & 150 & 5 & s-anomaly & 0.08(28) & \cite{Persson2020} \\
& 151 & 5/2 & 152 & 3 & s-anomaly & 0.50(6) & \cite{Persson2020} \\
& 151 & 5/2 & 153 & 5/2 & s-anomaly & -0.64(3) & \cite{Brand1981}\\
\\
Gd & 155 & 3/2 & 157 & 3/2 & s-anomaly & 0.106(24) & \cite{Persson2018GdNd} \\
\\
Dy & 161 & 5/2 & 163 & 5/2 & 4f$^{10}$6s6p $^{5}$K$_{8}$ & 0.019(16) & \cite{CLARK1982}
\\
&  &&&  & 4f$^{10}$6s6p $^{5}$K$_{9}$ & 0.025(11) & \cite{CLARK1982} \\
&  &&&  & 4f$^{10}$6s6p $^{5}$I$_{8}$ & -0.116(19) & \cite{CLARK1982} \\
&  &&&  & 4f$^{10}$6s6p $^{5}$H$_{7}$ & -0.176(36) & \cite{CLARK1982} \\
\\
Yb & 171 & 1/2 & 173 & 5/2 & 6s6p $^{3}$P$_{1}$ & -0.386(5) & \cite{BUDICK1970} \\
   &     &     &     &     &                    & -0.3857(51)& \cite{Atkinson2019}\\
&  &&&  & 4f$^{13}$5d6s$^{2}$ $^{3}$P$_{1}$ & 0.066(22) & \cite{BUDICK1970} \\
&  &&&  & Yb$^{+}$ 6s $^{2}$S$_{1/2}$, s-anomaly & -0.425(5) & \cite{Fisk1997}%
\cite{Munch1987} \\
\\

Lu & 175 & 7/2 & 176 & 7 & 5d6s$^{2}$ $^{2}$D$_{3/2}$ & 0.02(15) & \cite{Brenner1985}
\\
&  &&&  & 5d6s$^{2}$ $^{2}$D$_{5/2}$ & 0.19(15) & \cite{Brenner1985} \\
&  &&&  & 5d6s6p $^{4}$P$_{1/2}$ & 0.40(24) & \cite{Witte2002} \\
&  &&&  & 5d6s6p $^{4}$P$_{3/2}$ & 1.62(25) & \cite{Witte2002} \\
&  &&&  & 5d6s6p $^{4}$P$_{5/2}$ & 0.0(27) & \cite{Witte2002} \\
&  &&&  & 5d6s6p $^{4}$F$_{3,5,7/2}$, s-anomaly & 0.48(8) & \cite{Kuhnert1983}%
, $\mu _{I}$ from \cite{Brenner1985} \\
&  &&&  & 6s$^{2}$8p $^{2}$P$_{1/2}$ & 1.84(90) & \cite{Witte2002} \\
&  &&&  & 6s$^{2}$8p $^{2}$P$_{3/2}$ & 0.55(22) & \cite{Witte2002} \\
& 175 & 5/2 & 177 & 7/2 & s-anomaly & -0.018(35) & \cite{Brenner1985} \\
& 176 & 7 & 176$^{m}$ & 1 & s-anomaly & 0.48(8) & \cite{Brenner1985} \\
\\

Re & 185 & 5/2 & 186 & 1 & 5d$^{5}$6s$^{2}$ $^{6}$S$_{5/2}$ & -1.36(17) & \cite{Armstrong1965}\cite{Buttgenbach1981} \\
& 185 & 5/2 & 187 & 5/2 & 5d$^{5}$6s$^{2}$ $^{6}$S$_{5/2}$ & 0.031(8) & \cite{Armstrong1965} \\
& 185 & 5/2 & 187 & 5/2 & s-anomaly & 0.027(5) & \cite{Burger1982} \\
& 185 & 5/2 & 188 & 1 & 5d$^{5}$6s$^{2}$ $^{6}$S$_{5/2}$ & -1.28(28) & \cite{Armstrong1965}\cite{Buttgenbach1981} \\
\\

Ir & 191 & 3/2 & 193 & 3/2 & s-anomaly & -0.64(7) & \cite{Buttgenbach1978} \\
\\
\\
Au & 197 & 3/2 & 177 & 11/2 & 6s $^{2}$S$_{1/2}$ - 6p $^{2}$p$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 7.7(8) & \cite{Barzakh2020}\\
& 197 & 3/2 & 177 & 11/2 & s-anomaly & 11.4(14) & \cite{Barzakh2020}\\
& 197 & 3/2 & 185 & 5/2 & s-anomaly & 9.4(30) & \cite{Barzakh2020}\\
& 197 & 3/2 & 186 & 3 & s-anomaly & 3.1(51) & \cite{Barzakh2020}\\
& 197 & 3/2 & 187 & 1/2 & s-anomaly & 12.7(84) & \cite{Barzakh2020}\\
& 197 & 3/2 & 189 & 1/2 & s-anomaly & 9.4(59) & \cite{Barzakh2020}\\

 & 197 & 3/2 & 189 & 11/2 & 6s $^{2}$S$_{1/2}$ - 6p $^{2}$p$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 6.0(10) & \cite{Barzakh2020}
\\
& 197 & 3/2 & 189 & 11/2 & s-anomaly & 8.6(16) & \cite{Barzakh2020}\\

& 197 & 3/2 & 191 & 3/2 & s-anomaly & -1.2(14) & \cite{Barzakh2020}\\
 & 197 & 3/2 & 191 & 11/2 & 6s $^{2}$S$_{1/2}$ - 6p $^{2}$p$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 7.9(8) & \cite{Barzakh2020}\\
& 197 & 3/2 & 191 & 11/2 & s-anomaly & 11.7(14) & \cite{Barzakh2020}\\

& 197 & 3/2 & 193 & 3/2 & s-anomaly & -0.5(11) & \cite{Barzakh2020}\\
 & 197 & 3/2 & 193 & 11/2 & 6s $^{2}$S$_{1/2}$ - 6p $^{2}$p$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 7.6(6) & \cite{Barzakh2020}\\
& 197 & 3/2 & 193 & 11/2 & s-anomaly & 11.2(11) & \cite{Barzakh2020}\\

& 197 & 3/2 & 194 & 1 & s-anomaly & 1.8(33) & \cite{Barzakh2020}\\
 & 197 & 3/2 & 195 & 11/2 & 6s $^{2}$S$_{1/2}$ - 6p $^{2}$p$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 7.5(8) & \cite{Barzakh2020}\\
& 197 & 3/2 & 195 & 11/2 & s-anomaly & 11.4(14) & \cite{Barzakh2020}\\

& 197 & 3/2 & 196 & 2& 6s $^{2}$S$_{1/2}$, s-anomaly & 8.69(26) & \cite{Schmelling1970} \cite{EKSTROM1980} \\
& 197 & 3/2 & 198 & 2 & 6s $^{2}$S$_{1/2}$, s-anomaly & 8.53(8) & \cite{Bout1967}
\cite{EKSTROM1980} \\
& 197 & 3/2 & 199 & 3/2 & 6s $^{2}$S$_{1/2}$, s-anomaly & 3.64(29) & \cite{Bout1967}
\cite{EKSTROM1980} \\
\\

Hg & 199 & 1/2 & 193 & 3/2 & 6s6p $^{3}$P$_{1}$ & -0.61(3) & \cite{Reimann1973} \\
& 199 & 1/2 & 193$^{m}$ & 13/2 & 6s6p $^{3}$P$_{1}$ & -1.0552(13) & \cite{Reimann1973} \\
& 199 & 1/2 & 195 & 1/2 & 6s6p $^{3}$P$_{1}$ & -0.1470(9) & \cite{Reimann1973} \\
& 199 & 1/2 & 195$^{m}$ & 113/2 & 6s6p $^{3}$P$_{1}$ & -1.038(3) & \cite{Reimann1973} \\
& 199 & 1/2 & 197 & 1/2 & 6s6p $^{3}$P$_{1}$ & -0.0778(7) & \cite{Reimann1973} \\
& 199 & 1/2 & 197$^{m}$ & 13/2 & 6s6p $^{3}$P$_{1}$ & -1.021(3) & \cite{Reimann1973} \\
& 199 & 1/2 & 199$^{m}$ & 13/2 & 6s6p $^{3}$P$_{1}$ & --0.960(9) & \cite{Reimann1973} \\
& 199 & 1/2 & 201 & 3/2 & 6s6p $^{3}$P$_{1}$ & -0.1467(6) & \cite{Reimann1973} \\
& 199 & 1/2 &&  & 6s6p $^{3}$P$_{2}$ & -0.15653(4) & \cite{Reimann1973} \\
& 199 & 1/2 &&  & Hg$^{+}$, 6s $^{2}$S$_{1/2}$, s-anomaly & -0.16257(5) & \cite{Burt2009} \\
& 199 & 1/2 & 203 & 5/2 & 6s6p $^{3}$P$_{1}$ & -0.796(16) & \cite{Reimann1973} \\
\\
Tl & 203 & 1/2 & 205 & 1/2 & 6p $^{2}$P$_{1/2}$ & 0.01035(15) & \cite{LURIO1956} \\
   &     &&&     & 6p $^{2}$P$_{3/2}$ & -0.16258(10) & \cite{GOULD1956} \\
   &     &&&     & 7s $^{2}$S$_{1/2}$ & 0.0294(81) & \cite{Chen2012} \\
   &     &&&     & 7p $^{2}$P$_{1/2}$ & 0.05(4) & \cite{Ranjit2014} \\

   & 205 & 1/2 & 187 & 9/2 & 6p $^{2}$P$_{1/2}$ - 7s $^{2}$s$_{1/2}$ ($\Delta _{p}-\Delta
_{s}$) & 2.02(97) & \cite{Barzakh2012}\\
   & 205 & 1/2 & 189 & 9/2 & 6p $^{2}$P$_{1/2}$ - 7s $^{2}$s$_{1/2}$ ($\Delta _{p}-\Delta
_{s}$) & 1.47(105) & \cite{Barzakh2012}\\
   & 205 & 1/2 & 191 & 9/2 & 6p $^{2}$P$_{1/2}$ - 7s $^{2}$s$_{1/2}$ ($\Delta _{p}-\Delta
_{s}$) & 1.56(60) & \cite{Barzakh2012}\\ 
   & 205 & 1/2 & 193 & 9/2 & 6p $^{2}$P$_{1/2}$ - 7s $^{2}$s$_{1/2}$ ($\Delta _{p}-\Delta
_{s}$) & 1.31(62) & \cite{Barzakh2012}\\

Pb & 207 & 1/2 & 191 & 13/2 & 6p7s $^{3}$P$_{1}$ - 6p$^2$ $^{1}$D$_{2}$  & -1.72(68) & \cite{Persson2018Pb} \\
   & 207 & 1/2 & 193 & 13/2 & 6p7s $^{3}$P$_{1}$ - 6p$^2$ $^{1}$D$_{2}$ & -1.86(58) & \cite{Persson2018Pb} \\
   & 207 & 1/2 & 195 & 13/2 & 6p7s $^{3}$P$_{1}$ - 6p$^2$ $^{1}$D$_{2}$ & -1.53(70) & \cite{Persson2018Pb} \\
   & 207 & 1/2 & 197$^{m}$ & 13/2 & 6p7s $^{3}$P$_{1}$ - 6p$^2$ $^{1}$D$_{2}$ & -1.68(123) & \cite{Persson2018Pb} \\

\\
Fr & 212 & 5 & 206$^{g}$ & 3 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & 0.026(30) & \cite{Zhang2015} \\
& 212 & 5 & 206$^{m}$ & 7 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & -0.058(27) & \cite{Zhang2015} \\
& 212 & 5 & 207 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & -0.349(29) & \cite{Zhang2015} \\
& 212 & 5 & 208 & 7 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & -0.014(46) & \cite{Zhang2015} \\
& 212 & 5 & 208 & 7 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & 0.032(38) & \cite{Grossman1999} \\
& 212 & 5 & 209 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & -0.368(29) & \cite{Zhang2015} \\
& 212 & 5 & 209 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & -0.339(31) & \cite{Grossman1999} \\
& 212 & 5 & 210 & 6 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & 0.009(32) & \cite{Zhang2015} \\
& 212 & 5 & 210 & 6 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & 0.007(28) & \cite{Grossman1999} \\
& 212 & 5 & 211 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta
_{s}-\Delta _{p}$) & -0.334(31) & \cite{Zhang2015} \\
& 212 & 5 & 211 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta _{s}-\Delta
_{p}$) & -0.331(34) & \cite{Grossman1999} \\
& 212 & 5 & 213 & 9/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & -0.328(34) & \cite{Zhang2015} \\
& 212 & 5 & 221 & 5/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & -0.704(42) & \cite{Zhang2015} \\
\\
Ra & 211 & 5/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & 0.6(2) & \cite{Ahmad1983} \\
   & 221 & 5/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & -0.3(8) & \cite{Ahmad1983} \\
   & 223 & 3/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & 0.6(5) & \cite{Ahmad1983} \\
   & 225 & 1/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & 0.4(3) & \cite{Ahmad1983} \\
   & 225 & 1/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ & 0.8(4) & \cite{Arnold1987} \\
   & 227 & 3/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & 0.3(4) & \cite{Ahmad1983} \\
   & 229 & 5/2 & 213 & 1/2 & 7s $^{2}$S$_{1/2}$ - 7p $^{2}$P$_{1/2}$ ($\Delta_{s}-\Delta _{p}$) & 0.6(4) & \cite{Ahmad1983} \\
\\
U & 233 & 5/2 & 235 & 7/2 & 5f$^{3}$6d7s$^{2}$ $^{5}$L$_{6}$ & 0.84(31) & \cite{Gangrsky1997} \\
&  &&&  & 5f$^{3}$6d7s7p $^{7}$M$_{7}$ & 1.32(31) & \cite{Gangrsky1997} \\
&  &&&  & 5f$^{3}$6d7s7p $^{7}$L$_{6}$ & 1.19(89) & \cite{Gangrsky1997} \\
\end{longtable}
\end{landscape}
\newpage

\renewcommand{\arraystretch}{1.0}

\footnotesize

\newpage
\bibliographystyle{plain} 

\bibliography{hfa-2023}

\end{document}